OpenStudy (anonymous):

- 1/2 + 1/3 A.) - 1/6 B.) - 1/3 C.) 2/3 D.) 1/6??????????????????????

OpenStudy (mertsj):

What is the common denominator?

OpenStudy (anonymous):

6

OpenStudy (mertsj):

Write both fractions with a denominator of 6

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

1+1 = 2 2 times #3 = 6

OpenStudy (mertsj):

\[-\frac{3}{6}+\frac{2}{6}\]

OpenStudy (anonymous):

5/6

OpenStudy (mertsj):

Is that what you got when you wrote the common denominator?

OpenStudy (anonymous):

:/

OpenStudy (mertsj):

What is -3 + 2 ???

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

- 1/3?

OpenStudy (mertsj):

Then what is -3 sixths + 2 sixths?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

-3 +2=-1

OpenStudy (mertsj):

Then what is - 3 sixths + 2 sixths?

OpenStudy (mertsj):

To add two fractions with the same denominator: 1. Add the numerators 2. Leave the denominator alone.

OpenStudy (anonymous):

umm...5/6

OpenStudy (mertsj):

Now do this problem: \[\frac{2}{7}+\frac{3}{7}\]

OpenStudy (anonymous):

5/7

OpenStudy (mertsj):

And this one: \[\frac{5}{8}+\frac{-2}{8}\]

OpenStudy (anonymous):

-3/8

OpenStudy (mertsj):

That would be +3/8 since 5-2=3

OpenStudy (mertsj):

Now this one: \[\frac{-1}{5}+\frac{4}{5}\]

OpenStudy (anonymous):

+3/5

OpenStudy (mertsj):

And now this one: \[\frac{-3}{6}+\frac{2}{6}\]

OpenStudy (anonymous):

+1/6

OpenStudy (mertsj):

Which is bigger 3 or 2?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

cues its a positive

OpenStudy (mertsj):

Would you rather have 3 dollars or 2 dollars?

OpenStudy (anonymous):

3 but theres a negitive sign

OpenStudy (mertsj):

Why would you rather have 3 dollars than 2 dollars?

OpenStudy (anonymous):

cues it's more

OpenStudy (mertsj):

So I ask again...which is bigger 3 or 2?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

what about the negitive signs?

OpenStudy (mertsj):

And in this problem, is the 3 positive or negative?

OpenStudy (anonymous):

ignore

OpenStudy (anonymous):

negative

OpenStudy (mertsj):

Therefore your answer should be negative.

OpenStudy (mertsj):

Now. Answer the problem you posted at the very beginning. What is the answer?

OpenStudy (anonymous):

-2/3 :/

OpenStudy (anonymous):

-1/3 cues the two go's into one?

OpenStudy (mertsj):